# Alexandre M. Pombo11footnotetext: Corresponding author.    Lorenzo Pizzuti    and Alessandra di Giacomo

### v1

@@ 4.1 Domain of existence
Before analysing the other extreme, $\omega\to\omega_{-}$, it is important to point out that $Q$-balls stability is typically assessed on three levels [tsumagari2009physics]. i) Classical (linear) stability: along a solution branch parametrised by the frequency $\omega$, configurations are linearly stable if $\frac{dQ}{d\omega}<0$; turning points where $\frac{dQ}{d\omega}=0$ signal the onset (or loss) of one unstable mode. ii) Absolute (quantum) stability against decay into free quanta: the state is absolutely stable if its energy per charge satisfies $E/Q<\mu$; if $E/Q>\mu$ the $Q$-ball is at best metastable to particle emission. iii) Fission (fragmentation) stability: stability against splitting into smaller $Q$-balls is ensured when the energy is sub-additive, i.e., $E(Q_{1}+Q_{2})\leqslant E(Q_{1})+E(Q_{2})$ for all decompositions; a practical diagnostic is that $E/Q$ decreases with $Q$ along the branch (so larger-$Q$ states do not lower energy by fission). These criteria are model-independent and apply equally to polynomial and axion-like potentials; the size of the $\omega$-window in which they are met can differ across models.
@@
Of interest for our analysis is then the region of absolute stability for both the polynomial and axion-like $Q$-ball solutions. Which we summarize to be $\omega_{\rm stab}\in[0.62,0.82]\,$ while for the axion-like is $\omega_{\rm stab}\in[0.42,\,0.78]$.
@@ 5 Building rotation curves of a galaxy-size $Q$-ball
When converting to astrophysical units, one simply reinstates the appropriate factor of $c^{2}$ through $\rho=\varepsilon/c^{2}$. When studying $Q$-balls as disk galaxies’ DM halos, we are mainly focused on quantities defined at the equatorial plane – where observational data are obtained. As such, it is convenient to switch from spherical coordinates $(r,\theta,\phi)$ – where $Q$-balls are numerically computed –, to cylindrical coordinates $(R,z,\phi)$. This transformation enables a direct analysis of galactic rotation curves based on the gravitational potential and velocity profiles in the $z=0$ plane. In cylindrical coordinates $(R,z,\varphi)$, the field ansatz becomes:
@@
where now we have used $\hat{X}$ to explicitly highlight a quantity written in code units. For a given frequency, $\omega$, the corresponding physical mass density, acting as the source term of the Newtonian gravitational potential, is recovered through:
@@
The quantities $(\mu,\rho_{0})$ are determined by matching the dimensionless solution to observational data (see Section 7). Specifically, $\mu$ sets the radial scale, while $\rho_{0}$ is proportional to the total halo ($Q$-ball) mass. From the solution of (5.6), the gravitational potential in the equatorial plane generated by an axially symmetric mass distribution is given by  [BinneyTremaine]:
@@ 7 Results
Among the 175 galaxies in the SPARC catalogue, we consider a subset of 52 galaxies whose observed rotation curves extend up to at least 20 kpc (with more than 10 data points), in order to capture the outer distribution of the DM halo described by the $Q$-ball model. The total rotation curve of a galaxy is then expressed by including the contribution of the $Q$-ball:
@@
$BIC=k\ln(n)-2\ln(\ell)\,$, (7.2)
where $k$ is the number of free parameters, $n$ the number of data points and $\ell$ the likelihood of the fit. A model $A$ is significantly preferred over a model $B$ if $\Delta BIC=BIC_{B}-BIC_{A}\gtrsim 6$. In this analysis, the number of parameters does not change from the polynomial to the axion-like model, so the $\Delta BIC$ reduces to the difference of the log likelihoods (i.e. $\Delta\chi^{2}$).
@@
In Fig. 5 we show, as an example, the best fit axion-like and polynomial $Q$-ball rotation curves (green dashed lines), the baryonic contribution (yellow curves) and the total contribution (baryonic + $Q$-ball) along with its 68% confidence region (orange solid lines and light blue shaded area) compared to the observed curve (blue points with errorbars) for the spiral galaxy NGC 4183 in the Ursa Major cloud.
@@
Figure 5: Example of fitted rotation curves for the galaxy NGC 4183 of the SPARC sample. The orange solid lines and the light blue region indicate the best fit model from (7.1) and the 68% confidence region. Blue points with error bars are the observed rotation curves. The green dashed lines represent the best-fit rotation curve contributions of the $Q$-ball, while the yellow lines refer to the combined contribution of the baryonic components. Left: axion-like potential, $U_{A}$. Right: polynomial potential, $U_{P}$.
For the chosen galaxy, the $Q$-ball model fits reasonably well the observed rotation curve, with a reduced $\chi^{2}_{\rm red}=0.35,0.26$ for the polynomial and axion-like case, respectively. The axion-like potential performs slightly better ($\Delta BIC=1.68$), which is a common trend found on all the galaxies in our sample. Overall, the axion-like model is mildly preferred for 33 objects, with 12 exhibiting $\Delta BIC>6$. Only three galaxies indicate a strong ($\Delta BIC>6$) preference for the polynomial case, and they are all characterised by a $\chi^{2}_{\rm red}>1$. In total, 18 (15) galaxies exhibit a $\chi^{2}_{\rm red}<1$ when fitting the axion-like (polynomial) $Q$-ball; the number increases to 35 for $\chi^{2}_{\rm red}<2$.
@@
It is important to remark here that we are adopting an illustrative, simplified model which neglects the gravitational term in the scalar field action and does not account for the complex physical mechanisms in the galactic environment. Thus, this comparison should be considered cautiously as a proof-of-concept for future astrophysical applications. However, the presence of a toroidal-shaped DM distribution provided by the spinning $Q$-ball seems to provide a natural description for the observed DM distribution of disk galaxies.
@@
Fig. 6 shows the estimated best fit log particle mass, log halo mass and frequency $\omega$ parameters with their 68% intervals. The red points and bars indicate fits with a $\chi^{2}_{\rm red}$ larger than 1.0. For both polynomial and axion-like potentials, we found that in the majority of the cases (39 galaxies) the particle mass is constrained to be between $28<\log_{10}(m/\text{eV})<-26.5$, which is consistent with the results presented in [Mourelle24, Mourelle25] and the expected maximum mass for ultra-light DM models [Marsh_2022, lopez2025scaling].
@@
$\text{log}_{10}(M_{\rm h}/\text{M}_{\odot})=0.47\pm 0.06\,,\qquad\text{log}_{10}(m/\text{eV})=-0.20\pm 0.04\,,\qquad\omega=0.81\pm 0.01\,$, (7.3)
@@
$\text{log}_{10}(M_{\rm h}/\text{M}_{\odot})=0.40\pm 0.06\,,\qquad\text{log}_{10}(m/\text{eV})=-0.20\pm 0.04\,,\qquad\omega=0.72\pm 0.2\,$, (7.4)
@@ 8 Conclusions
In this paper, we examined whether dynamically robust, rotating scalar $Q$-balls can reproduce the DM contribution of disk-galaxy rotation curves. We presented a theoretical and numerical framework for two classes of self-interactions: the standard sextic polynomial potential, $U_{P}$, and an axion–like periodic potential, $U_{A}$, that reduces to the polynomial form at small amplitudes ($\omega\to\mu)$, but departs from it at high (non-linear regime).
@@
To compute these solutions efficiently and in an approximated closed form, we presented for the first time the SpectralPINN, an in-house developed hybrid solver that embeds pseudo-spectral basis functions (PSM) directly inside a PINN optimisation loop. The result is a solver that preserves the exponential accuracy of spectral approaches while retaining the flexibility of neural networks residual minimisation and yields compact representations of both spherically symmetric and spinning $Q$-balls for the two potentials considered.
@@
The computed solutions, and their domain of existence, are analysed for the first time in the axion-like case. For the spinning and spherical configurations, both models coincide at $\omega\to\mu$ as expected from the low field approximation. On the other hand, toward $\omega\to\omega_{-}$ the polynomial $U_{P}$ is monotonicaly increasing (after reaching a minimum for the spinning case) ($E,\,Q\to+\infty$), while the axion-like solution reaches a local (global) maximum of $E$ ($Q$) (once again after a reaching a minimum for the spinning case), turns over and continues to a high-energy/low-$Q$ branch with $E/Q>\mu$. Absolute stability ($E/Q<\mu$) occurs only in an intermediate band – narrow for $U_{P}$, broader for $U_{A}$ – while classical stability tracks segments with $dQ/d\omega<0$.
@@
An empirical assessment on a quality–controlled subset of 52 galaxies from the SPARC catalog [lelli_2016_sparc] shows that both potentials can fit observed rotation curves competitively with standard halo prescriptions. The axion–like potential is mildly preferred overall, delivering the lowest BIC in 33 systems (with a dozen strong preferences) and achieving $\chi^{2}_{\rm red}<1$ in more cases. In most galaxies the inferred scalar–particle mass clusters around $m_{\phi}\sim 10^{-27}\,\mathrm{eV}$, in line with expectations for solitonic DM scenarios on galactic scales (e.g. [Marsh_2022, Mourelle24]). A small number of objects favour unrealistically low masses and excessively large halo normalisations; however, these can be considered as artifacts of our simplified setup and of residual systematics in individual rotation curves, rather than viable astrophysical solutions.
@@
Note that the present work is intentionally proof–of–concept: we neglected the field’s self–gravity in the equations of motion, and adopted a simplified treatment for the galaxy environment which does not account for all observational systematics. These choices isolate the core question of viability but limit the completeness of the inference. The natural extension of this analysis will incorporate self–gravity in the action and confront the model jointly with complementary probes such as stellar/gas kinematics, lensing, and the baryonic Tully–Fisher relation. Moreover, the inclusion of dwarf galaxies’ rotation curves in the sample may provide additional clues on the ultra-light DM and fuzzy DM hypothesis, with many studies suggesting a larger particle mass $m>10^{-23}$eV (e.g. [Zoutendijk21, Banares23]).
@@
Taken together, our results identify rotating scalar $Q$-balls, especially with axion–like self–interactions, as credible solitonic candidates for the dark components of disk galaxies at the level of their rotation curves, and they establish SpectralPINN as a practical tool for generating accurate, storage–light solutions amenable to astrophysical forward modeling. With fuller dynamics and a broader data confrontation, this framework can sharpen both astrophysical and particle–physics interpretations of DM at galaxy scales.

### v2

@@ 4.1 Domain of existence
Before analysing the other extreme, $\omega\to\omega_{-}$, it is important to point out that $Q$-balls stability is typically assessed on three levels [52]: i) Classical (linear) stability, along a solution branch parametrised by the frequency $\omega$, configurations are linearly stable if $\frac{dQ}{d\omega}<0$ – turning points where $\frac{dQ}{d\omega}=0$ signal the onset (or loss) of one unstable mode; ii) absolute (quantum) stability against decay into free quanta – the state is absolutely stable if its energy per charge satisfies $E/Q<\mu$, if $E/Q>\mu$ the $Q$-ball is at best metastable to particle emission; iii) fission (fragmentation) stability: stability against splitting into smaller $Q$-balls is ensured when the energy is sub-additive, i.e., $E(Q_{1}+Q_{2})\leqslant E(Q_{1})+E(Q_{2})$ for all decompositions – a practical diagnostic is that $E/Q$ decreases with $Q$ along the branch (so larger-$Q$ states do not lower energy by fission). These criteria are model-independent and apply equally to polynomial and axion-like potentials; the size of the $\omega$-window in which they are met can differ across models.
@@
Of interest for our analysis is then the region of absolute stability for both the polynomial and axion-like $Q$-ball solutions. Which we summarize to be $\omega_{\rm stab}\in[0.62,0.82]$ for the polynomial, while for the axion-like is $\omega_{\rm stab}\in[0.42,\,0.78]$.
@@ 5 Building rotation curves of a galaxy-size $Q$-ball
When converting to astrophysical units, one simply reinstates the appropriate factor of $c^{2}$ through $\rho=\varepsilon/c^{2}$. When studying $Q$-balls as disk galaxies’ DM halos, we are mainly focused on quantities defined at the equatorial plane – where observational data are obtained. As such, it is convenient to switch from spherical coordinates $(r,\theta,\varphi)$ – where $Q$-balls are numerically computed –, to cylindrical coordinates $(R,z,\varphi)$. This transformation enables a direct analysis of galactic rotation curves based on the gravitational potential and velocity profiles in the $z=0$ plane. In cylindrical coordinates $(R,z,\varphi)$, the field ansatz becomes:
@@
where now we have used $\hat{X}$ to explicitly highlight a quantity written in code units. For a given frequency, $\omega/\mu$, the corresponding physical mass density, acting as the source term of the Newtonian gravitational potential, is recovered through:
@@
The quantities $(\mu,\rho_{0})$ are determined by matching the dimensionless solution to observational data (see Section 7). Specifically, $\mu$ sets the radial scale, while $\rho_{0}$ is proportional to the total halo ($Q$-ball) mass. From the solution of (5.6), the gravitational potential in the equatorial plane generated by an axially symmetric mass distribution is given by  [3]:
@@ 7 Results
Among the 175 galaxies in the SPARC catalogue, in order to minimise selection effects, we consider disk galaxies with at least ten measured rotation-curve points, resulting in a final sample of 116 objects spanning a wide range of masses and morphologies. This choice allowed us to assess the phenomenological signature of the $Q$-ball model across a variety of disk galaxy population, rather than focusing on a restricted subsample optimised for a specific halo profile.
The total rotation curve of a galaxy is then expressed by including the contribution of the $Q$-ball:
@@
$\text{BIC}=k\ln(n)-2\ln(\ell)\ ,$ (7.2)
[…]
To evaluate the viability of the $Q$-balls, we further compare both the polynomial and the axion-like models with a fit obtained assuming standard Navarro-Frenk-White (NFW) halo profile [34] for the DM distribution. The latter’s associated circular velocity at radius $r$ reads:
$v^{2}_{NFW}(r)=\frac{GM_{NFW}(r;r_{200},r_{s})}{r}\ ,$ (7.3)
where
$M(r;r_{200},r_{s})=M_{200}\frac{\ln(1+r/r_{s})-(r/r_{s})/(1+r/r_{s})}{\ln(1+c_{200})-c_{200}/(1+c_{200})}\ .$ (7.4)
With $r_{s}$ and $r_{200}$ are the two free parameters characterising the halo, $c_{200}=r_{200}/r_{s}$ and $M_{200}=(100H^{2}(z)/G)\times r_{200}^{3}$. Finally, $H(z)$ is the Hubble parameter at redshift $z$.
@@
In Fig. 5 we show, as an example, the best fit axion-like and polynomial $Q$-ball rotation curves (green dashed lines in the left and central plots), the baryonic contribution (yellow curves) and the total contribution (baryonic + $Q$-ball) along with its 68% confidence region (orange solid lines and light blue shaded area) compared to the observed curve (blue points with error bars) for the spiral galaxy NGC 4183 in the Ursa Major cloud.
@@
Figure 5: Example of fitted rotation curves for the galaxy NGC 4183 of the SPARC sample. The orange solid lines and the light blue region indicate the best fit model from (7.1) and the 68% confidence region. Blue points with error bars are the observed rotation curves. The green dashed lines represent the best-fit rotation curve contributions of the $Q$-ball (left and central plots) and of the NFW DM profile (right), while the yellow lines refer to the combined contribution of the baryonic components. Left: axion-like potential, $U_{A}$. Central: polynomial potential, $U_{P}$. Right: standard NFW DM distribution.
[…]
@@
It is important to remark here that we are adopting an illustrative, simplified model which neglects the gravitational term in the scalar field action and does not account for the complex physical mechanisms in the galactic environment. Thus, this comparison should be considered cautiously as a proof-of-concept for future astrophysical applications. However, the toroidal morphology characteristic of spinning $Q$-balls seems to translate, in the equatorial plane, into a distinctive family of circular–velocity profiles $v(R)$ that can reproduce the dark component of disk–galaxy rotation curves.
@@
Fig. 7 shows the estimated best fit log particle mass, log halo mass and frequency $\omega$ parameters with their 68% intervals. The red points and bars indicate fits with a $\chi^{2}_{\rm red}$ larger than 1.0. For both polynomial and axion-like potentials, we found that in the majority of the cases (39 galaxies) the particle mass is constrained to be between $28<\log_{10}(m/\text{eV})<-26.5$, which is consistent with the results presented in [32, 33] and the expected maximum mass for ultra-light DM models [28, 26].
@@
$\text{log}_{10}(M_{\rm h}/\text{M}_{\odot})=12.65\pm 0.07\,,\qquad\text{log}_{10}(m/\text{eV})=-27.35\pm 0.04\,,\qquad\omega=0.81\pm 0.01\,$, (7.5)
@@
$\text{log}_{10}(M_{\rm h}/\text{M}_{\odot})=12.44\pm 0.06\,,\qquad\text{log}_{10}(m/\text{eV})=-27.15\pm 0.04\,,\qquad\omega=0.69\pm 0.02\,$, (7.6)
@@ 8 Conclusions
In this paper, we examined whether dynamically robust, rotating scalar $Q$-balls can reproduce the DM contribution to disk-galaxy rotation curves. We presented a theoretical and numerical framework for two classes of self-interactions: the standard sextic polynomial potential, $U_{P}$, and an axion-like periodic potential, $U_{A}$, that reduces to the polynomial form at small amplitudes ($\omega\to\mu)$, but departs from it at high (non-linear regime).
@@
To compute these solutions efficiently and in an approximated closed form, we presented for the first time the SpectralPINN, an in-house developed hybrid solver that embeds pseudo-spectral basis functions directly inside a PINN optimisation loop. The result is a solver that preserves the exponential accuracy of spectral approaches while retaining the flexibility of neural networks residual minimisation and yields compact representations of both spherically symmetric and spinning $Q$-balls for the two potentials considered.
@@
The computed solutions, and their domain of existence, are analysed for the first time in the axion-like case. For the spinning and spherical configurations, both models coincide at $\omega\to\mu$ as expected from the low field approximation. On the other hand, toward $\omega\to\omega_{-}$ the polynomial $U_{P}$ is monotonically increasing (after reaching a minimum for the spinning case) ($E,\,Q\to+\infty$), while the axion-like solution reaches a local (global) maximum of $E$ ($Q$) (once again after reaching a minimum for the spinning case), turns over and continues to a high-energy/low-$Q$ branch with $E/Q>\mu$. Absolute stability ($E/Q<\mu$) occurs only in an intermediate band – narrow for $U_{P}$, broader for $U_{A}$ – while classical stability tracks segments with $dQ/d\omega<0$.
An empirical assessment on a quality-controlled subset of 116 galaxies from the SPARC catalog [24] shows that both potentials can fit observed rotation curves with statistically acceptable quality in many systems, and can be competitive with the standard NFW DM model in a subset of galaxies. More in detail, the axion-like potential is mildly preferred over the polynomial, delivering the lowest BIC in 22 systems, while the polynomial is preferred only for 10 galaxies. While the NFW provides the lowest BIC for 84 systems, the preference is strong only for a small ($\lesssim 10\%$) fraction of the sample.
@@
In most galaxies the inferred scalar-particle mass clusters around $m_{\phi}\sim 10^{-27}\,\mathrm{eV}$, in line with expectations for solitonic DM scenarios on galactic scales (e.g. [28, 32]). Around 30% of the objects analysed favour unrealistically low masses and excessively large halo normalisations; however, these can be considered as artifacts of our simplified setup and of residual systematics in individual rotation curves, rather than viable astrophysical solutions.
We emphasize that the present work is intentionally proof-of-concept. By construction, we neglect the scalar field self–gravity in the equations of motion and adopt a simplified description of the galactic environment, which does not account for all observational systematics. These choices isolate the core question of phenomenological viability, but naturally limit the completeness of the inference and imply that we do not expect to outperform empirical halo profiles such as NFW across the full galaxy sample.
@@
The natural prosecution of this analysis will incorporate self-gravity directly in the action and confront the model jointly with complementary probes, including stellar and gas kinematics, gravitational lensing, and the baryonic Tully-Fisher relation. Moreover, extending the sample to include dwarf-galaxy rotation curves may provide additional insight into the ultra-light and fuzzy DM regime, with several studies suggesting a larger particle mass, $m\gtrsim 10^{-23},\mathrm{eV}$ (e.g. [55, 2]).
@@
Taken together, our results identify rotating scalar $Q$-balls, especially with axion-like self-interactions, as credible solitonic candidates for the dark components of disk galaxies at the level of their rotation curves, and they establish SpectralPINN as a practical tool for generating accurate, storage-light solutions amenable to astrophysical forward modeling. With fuller dynamics and a broader data confrontation, this framework can sharpen both astrophysical and particle-physics interpretations of DM at galaxy scales.